OpenStudy (anonymous):

Solve:

OpenStudy (anonymous):

\[(x+3)^{2} + 2x = -3\]

OpenStudy (anonymous):

itll be -2

OpenStudy (eyust707):

we can expand the ( )'s \[(x+3) (x+3) + 2x = -3\]then foil out the ( )'s\[x^2 + 6x + 9 + 2x = -3\]\[x^2 + 8x + 9 = -3\]add 3 to each side so our quadratic is equal to 0\[x^2 +8x + 12 = 0\]factor\[(x + 6) (x+ 2)\]\[x=-6\] and \[x= -2\]

OpenStudy (anonymous):

Thank you...I calculated the parenthesis first and it screwed everything up. Thanks!